OpenStudy (anonymous):

Help Me .. I Need Help on this Problem Solve for x. 12 – 3x – x = 8 A. –10 B. –5 C. 1 D. 2

OpenStudy (phi):

12 -3x - x = 8 what is -3 x's take away another x ?

OpenStudy (phi):

you should get -3x - x = -4x and your problem is 12 -4x = 8 you could add -12 to both sides of the equation 12 - 12 -4x = 8 -12 now simplify both sides. what do you get ?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

abcdefg . lol

OpenStudy (anonymous):

12-3x-x=8 12-4x=8 -4x=8-12 -4x=-4 x=-4/-4 x=1 so the answer is C: 1

OpenStudy (anonymous):

did u understand???

OpenStudy (anonymous):

Hell Yeah . Thanks :)

OpenStudy (anonymous):

ur welcome :)

OpenStudy (anonymous):

LOL . Thnks @khadeeja Can You help me on this Problem Right Here I Need Help on this Problem ??? ~HELP !! FAST !! I NEED HELP ON THIS !! ♥ ↓ ----------------------↓------------↓ Identify the solution set of the given inequality using the given replacement set. x < –5; {–9, –6.5, –5.2, –5, 0, 4.2} A. {–9, –6.5, –5.2} B. {–9, –6.5, –5.2, –5} C. {–5, 0, 4.2} D. {0, 4.2}